x=1+the square root of x squared + x + 2x times the square root of x - the square root of x + 1 + 2 times the square root of x
Answer by AnlytcPhil(1806)

x=1+the square root of x squared + x + 2x times the 
square root of x - the square root of x + 1 + 2 times 
the square root of x
         ____      _    ___     _ 
x = 1 + Öx2+x + 2xÖx - Öx+1 + 2Öx

Check to see if 0 is a solution

         ____        _    ___     _ 
0 = 1 + Ö02+0 + 2(0)Ö0 - Ö0+1 + 2Ö0

0 = 1 + 0 + 0 - 1 + 0

0 = 0

Yup, sure is. That's the solution. So x=0

Edwin
AnlytcPhil@aol.com